one conveyor belt requires 20 minutes to fill a large truck with ore. A second conveyor belt requires 30 minutes to fill the same truck with ore. how long would it take if both conveyor belts were working together to fill the truck with ore?
Answer by checkley79(3341)

1/20+1/30=1/x
(30+20)/20*30=1/x
50/600=1/x
50x=600
x=600/50
x=12 hours to fill the truck usung both conveyers.